# footnotetext: ${}^{\text{\Letter}}$ Corresponding Authors

### v1

@@ Appendix B Negativity and Positivity of Enthalpy for the Arbitrary Dimensional FRW Universe
In Einstein gravity, use the Eqs.(15), (23) and (18), we obtain the enthalpy of $n$-dimensional FRW universe is
@@
In Einstein-Gauss-Bonnet gravity, use the Eqs.(34), (39) and (18) we obtain the enthalpy of $n$-dimensional FRW universe is then
@@
    $R_{A}=\sqrt{-2\widetilde{\alpha}}\,.$ (71)
    It corresponds to a thermodynamic singularity.
Appendix C Constraints on the Perfect Fluid of the FRW Universe from Inversion Point ($\mu=0$)
In this appendix, we give the constraints on the perfect fluid of the FRW universe from the inversion point ($\mu=0$).
In the scenario where $\mathscr{H}<0$, it is intriguing to observe the presence of an inversion point, which corresponds to a peak in the $T$-$P$ diagram during the Joule-Thomson expansion. Exploring the deeper physical significance of this phenomenon is essential. Simplified discussions are presented below.
Equating the Hawking temperature (7) and the inversion temperature (53), we obtain
$\frac{\dot{R}_{A}}{HR_{A}}=\frac{2}{n-2}\,,$ (72)
which together with (5) results to the following relation
$\left(\dot{H}-\frac{k}{a^{2}}\right)R_{A}^{2}=-\frac{2}{n-2}\,,$ (73)
and by using (4), we finally obtain [^8] When $n=4$, the expression (74) reduces to $\ddot{a}=0$ at the inversion point, which indicating a transition between acceleration ($\ddot{a}>0$) and deceleration ($\ddot{a}<0$), or vice versa Abdusattar:2021wfv .
$\left(\dot{H}-\frac{k}{a^{2}}\right)+\frac{2}{n-2}\left(H^{2}+\frac{k}{a^{2}}\right)=0\,.$ (74)
The above condition (74) also corresponds to a constraint on the perfect fluid, which is still obtained by using equations (4), (5), (11), and the constraint is [^9] It is also consistent with the result of constraint on the perfect fluid from negative surface gravity obtained in (60).
$1+\frac{p}{\rho}=\frac{4}{(n-1)(n-2)}\,.$ (75)
These observations highlights the noticeable influence of spacetime dimensions on the perfect fluid and the evolution of the FRW universe.

### v2

@@ Appendix B Negativity and Positivity of Enthalpy for the Arbitrary Dimensional FRW Universe
Enthalpy is an important thermodynamic quantity that includes internal energy and the pressure-volume work to push surrounding systems. It has now become an effective tool for studying the thermodynamic behavior of gravitational systems, including the standard FRW spacetime. This section will analyze the positivity and negativity of enthalpy in arbitrary dimensional FRW universes, with the core being to find the general conditions that determine its sign.
In Einstein gravity, using Eqs.(15), (23) and (18), we obtain the enthalpy of the $n$-dimensional FRW universe as
@@
In Einstein-Gauss-Bonnet gravity, using Eqs.(34), (39) and (18), we obtain the enthalpy of the $n$-dimensional FRW universe
@@
    $R_{A}=\sqrt{-2\widetilde{\alpha}}\,$, (79)
    which corresponds to a thermodynamic singularity.